An electric current is passed through a circuit containing two wires of same material connected in parallel. If the lenghts and radii of the wires are in the ratio 4/3 and 2/3 then the ratio of the currents passing through the wire will be?

Grade:11

1 Answers

Khimraj
7 years ago
Ohm`s law V = iR and for parallel combination V is same so I is inversely proportional to R And R = pl/A where l is length, A is cross section area and p is resistivity of wire. since both wire are of same material p is same soi1/i2 = R2/R1 = (l2/l1)*(r1/R2)² =(3/4)(2/3)² = 1/3